Solve by the substitution method.
1/3x+1/5y=-4/5
1/3x+4y=3

Answer by Alan3354(69443)

1/3x+1/5y=-4/5
1/3x+4y=3
============
1/3x = -1/5y - 4/5
1/3x = -4y + 3
-------------------
Sub for (1/3)x
-1/5y - 4/5 = -4y + 3
Multiply by 5
-y - 4 = -20y + 15
etc
-------------
Elimination is the best method for these 2 equations.